# A note on embracing exchange sequences in oriented matroids

### v1

@@ 1 Introduction
A natural relaxation of this problem arises when there are only $k<d+1$ point sets, and the goal is to find a colourful set containing, for each $i\in[k]$, $d_{i}$ points from the $i$-th set. When $k=2$ and the target sizes $(d_{1},d_{2})$ satisfy $d_{1}+d_{2}=d+1$, the problem admits a polynomial-time solution, as shown by Meunier, Mulzer, Sarrabezolles, and Stein [7]. In an effort to better understand this 2-colour case and to explore the possibility of extending polynomial-time algorithms to higher numbers of colours, Caoduro, Khodamoradi, Shepherd, and Paat [3] raised the following problem: Given two 0-embracing sets $A,B\subseteq\mathbb{R}^{d}$, that is, sets whose convex hulls contain the origin, is it always possible to transform $A$ into $B$ by a polynomial-length sequence of exchanges, where each step replaces a point of the current set with a point from $A\cup B$ and every intermediate set remains 0-embracing? We call the minimum number of such exchanges the 0-embracing exchange distance of $A$ and $B$.
Let us make a few remarks. First, it is not difficult to see that $A$ and $B$ may be assumed to consist of the vertices of two simplices. Second, it is worth noting that if we drop the requirement on the length of the sequence, then the existence of a suitable sequence is always guaranteed. To see this, consider the polytope $P=\{x\in\mathbb{R}^{2(d+1)}\colon 0\leq x\leq 1,\ [C_{A},C_{B}]x=0\}$, where $C_{A}$ and $C_{B}$ are the matrices whose columns are the points of $A$ and $B$, respectively. Since $A$ and $B$ are 0-embracing, there exist nonzero vectors $x_{A},x_{B}\in\mathbb{R}^{d+1}$ with minimal support such that $C_{A}x_{A}=0$ and $C_{B}x_{B}=0$. By scaling these vectors so that their maximum component is $1$, we uniquely determine their nonzero entries. Consequently, the vectors $x_{A},0]$ and $0,x_{B}]$ are vertices of $P$. Let $G$ denote the $1$-skeleton of $P$. Since the graph of a convex polytope is connected, there exists a path in $G$ from $x_{A},0]$ to $0,x_{B}]$. Each vertex of $P$ corresponds to a minimal 0-embracing set defined by the support of the vertex. Moving along an edge of $P$ corresponds to a pivot step (or basis exchange), wherein one index leaves the support and another enters. Thus, following the path in $G$ yields a sequence of exchanges where, at each step, a point is removed from the current minimal support and replaced by a point from $A\cup B$, maintaining the origin in the convex hull.
Based on computational experiments, Caoduro, Khodamoradi, Shepherd, and Paat [3] proposed the following conjecture.
Conjecture 1.1 (Caoduro, Khodamoradi, Shepherd, Paat).
@@
The conjecture remains open. The problem, however, resembles several long-standing conjectures concerning the symmetric exchange distance of basis-pairs in matroids, such as those proposed by Gabow [6], White [9], and Hamidoune [5]; see Section 3.2 for details. This observation naturally suggests that the problem may, in fact, be a special case of a more general, abstract question concerning matroids. The present note aims to formally state and popularize such a conjecture, and to initiate its study.
2 Notation
@@
3 Matroidal extension
Any finite set $E\subseteq\mathbb{R}^{d}$ defines an affine matroid $M=(E,\mathcal{C})$, where a subset $X\subseteq E$ is independent if its vectors are affinely independent. A minimal dependent set in $M$ is called a circuit. Each circuit can be assigned a sign pattern based on a nontrivial linear combination of its elements that sums to zero, resulting in a signed circuit. Reversing all signs yields another valid signed circuit, so each circuit gives rise to a pair of oppositely signed circuits. In this way, one obtains an oriented matroid from the underlying affine matroid. If $E$ is full-dimensional, then the bases of $M$ correspond to subsets of $E$ defining full-dimensional simplices; note that the rank of $M$ is $d+1$. A basis contains the origin precisely when, in the signed fundamental circuit of $\mathbf{0}$ oriented so that $\mathbf{0}$ has negative sign, every other element is positive. Equivalently, the vertices of the corresponding simplex then form a $\mathbf{0}$-embracing basis. Therefore, Conjecture 1.1 asserts that the $\mathbf{0}$-embracing exchange distance between any two $e$-embracing bases of the oriented affine matroid is at most the dimension plus one. As an extension to general oriented matroids, we propose the following.
@@ 3 Matroidal extension
In contrast to Conjecture 1.1, a key difference in the abstract oriented matroid setting is that it remains open whether any exchange sequence between two bases meeting the conjecture’s conditions exists at all, regardless of its length. This is in sharp contrast to the unoriented setting: if $A$ and $B$ are bases of an unoriented matroid $M$, the exchange axiom guarantees the existence of a sequence of exchanges transforming $A$ into $B$, and the shortest such sequence has length $|A\setminus B|$. No analogous statement holds in the oriented setting, not even for the oriented graphic matroid of directed graphs; see Example 3.2.
The main goal of the present note is to popularize this question as a possible research direction. As a first step towards the general setting, we verify that the conjecture holds when $M$ is the oriented graphic matroid of a directed graph (Theorem 3.3). In fact, we prove a slightly stronger statement: there exists an exchange sequence of length at most $r$ which is monotone in the sense that the intersection with the target basis does not decrease at every step. We also highlight the connection between our problem and certain conjectures on exchange distances of basis-pairs, and show that no analogous strengthening of the conjecture is possible (Example 3.6).
3.1 Directed graphs
@@
3.2 Matroid exchange analogues
@@
There is a clear similarity between Conjectures 3.1 and 3.5, which raises the question of whether the former can be strengthened by allowing symmetric exchanges. More formally, given two $e$-embracing bases $A$ and $B$ of a rank-$r$ oriented matroid, one may ask whether there always exists a sequence of at most $r$ symmetric exchanges transforming $(A,B)$ into $(B,A)$ such that every intermediate pair in the sequence consists of $e$-embracing bases. We show that such a strengthening is not possible, even when the oriented matroid in question is an affine matroid and the bases are disjoint.
@@ Acknowledgement.
Kristóf Bérczi is grateful to Marco Caoduro for introducing him to the original problem on 0-embracing sets, and to András Imolay, Benjamin Schröter and Lorenzo Vecchi for initial discussions on the matroidal extension. The research was supported by the Lendület Programme of the Hungarian Academy of Sciences – grant number LP2021-1/2021, by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund – grant numbers ADVANCED 150556 and ELTE TKP 2021-NKTA-62, and by Dynasnet European Research Council Synergy project – grant number ERC-2018-SYG 810115.

### v2

@@ 1 Introduction
A natural relaxation of this problem arises when there are only $k<d+1$ point sets, and the goal is to find a colourful set containing, for each $i\in[k]$, $d_{i}$ points from the $i$-th set. When $k=2$ and the target sizes $(d_{1},d_{2})$ satisfy $d_{1}+d_{2}=d+1$, the problem admits a polynomial-time solution, as shown by Meunier, Mulzer, Sarrabezolles, and Stein [9]. In an effort to better understand this 2-colour case and to explore the possibility of extending polynomial-time algorithms to higher numbers of colours, Caoduro, Khodamoradi, Paat, and Shepherd [4] raised the following problem: Given two 0-embracing sets $A,B\subseteq\mathbb{R}^{d}$, that is, sets whose convex hulls contain the origin, is it always possible to transform $A$ into $B$ by a polynomial-length sequence of exchanges, where each step replaces a point of the current set with a point from $A\cup B$ and every intermediate set remains 0-embracing? We call the minimum number of such exchanges the 0-embracing exchange distance of $A$ and $B$.
To see this, form a $(d+1)\times(2d+2)$ matrix $M$ whose columns are the coordinates of the points in $A\cup B$, augmented by an additional row consisting entirely of $1$’s. Let $b\in\mathbb{R}^{d+1}$ be the vector whose first $d$ entries are 0 and whose last entry is $1$. Each exchange of a point in $A$ with a point in $B$ can then be interpreted as traversing an edge of the polyhedron $P=\{x\in\mathbb{R}^{2d+2}\mid Mx=b,;x\geq 0\}$. Therefore, if $G$ denotes the $1$-skeleton of $P$, then following a path in $G$ yields a sequence of exchanges in which, at each step, one point is removed from the current simplex and replaced by another point from $A\cup B$, while preserving the property that the origin lies in the convex hull. Based on computational experiments, Caoduro, Khodamoradi, Paat, and Shepherd [4] proposed the following conjecture.
Conjecture 1.1 (Caoduro, Khodamoradi, Paat, Shepherd).
@@
Although the Hirsch conjecture is now known to be false in general, Conjecture 1.1 would follow from its original formulation (see [7]), since the polyhedron $P$ has $2d+2$ facets and dimension 2d+2(d+1)=d+12d+2-(d+1)=d+1. Consequently, the Hirsch bound would guarantee the existence of a path of length at most $d+1$ between the corresponding vertices in the $1$-skeleton of $P$.
Conjecture 1.1 remains open. The problem, however, resembles several long-standing conjectures concerning the symmetric exchange distance of basis-pairs in matroids, such as those proposed by Gabow [8], White [11], and Hamidoune [6]; see Section 3 for details. This observation naturally suggests that the problem may, in fact, be a special case of a more general, abstract question concerning matroids.
1.1 Basic notation
@@
1.2 Matroidal extension
Any finite set $E\subseteq\mathbb{R}^{d}$ defines an affine matroid $M=(E,\mathcal{C})$, where a subset $X\subseteq E$ is independent if its vectors are affinely independent. A minimal dependent set in $M$ is called a circuit. Each circuit can be assigned a sign pattern based on a nontrivial linear combination of its elements that sums to zero, resulting in a signed circuit. Reversing all signs yields another valid signed circuit, so each circuit gives rise to a pair of oppositely signed circuits. In this way, one obtains an oriented matroid from the underlying affine matroid. If $E$ is full-dimensional, then the bases of $M$ correspond to subsets of $E$ defining simplices; note that the rank of $M$ is $d+1$. A basis contains the origin precisely when, in the signed fundamental circuit of $\mathbf{0}$ oriented so that $\mathbf{0}$ has negative sign, every other element is positive. Equivalently, the vertices of the corresponding simplex then form a $\mathbf{0}$-embracing basis. Therefore, Conjecture 1.1 asserts that the $\mathbf{0}$-embracing exchange distance between any two $\mathbf{0}$-embracing bases of the oriented affine matroid is at most the dimension plus one. As an extension to general oriented matroids, we propose the following.
@@ 1.2 Matroidal extension
Closely related problems have been considered in the context of bounding the diameter of the cocircuit graph of an oriented matroid; see [1]. That work studies linear and polynomial diameter bounds, and establishes positive results for oriented matroids of low rank and corank, as well as for small ground sets via computer verification. In contrast to Conjecture 1.1, a key difference in the abstract oriented matroid setting is that it remains open whether any exchange sequence between two bases meeting the conjecture’s conditions exists at all, regardless of its length. This is in sharp contrast to the unoriented setting: if $A$ and $B$ are bases of an unoriented matroid $M$, the exchange axiom guarantees the existence of a sequence of exchanges transforming $A$ into $B$, and the shortest such sequence has length $|A\setminus B|$. No analogous statement holds in the oriented setting, not even for the oriented graphic matroid of directed graphs; see Example 2.1.
As a first step towards the general setting, we verify that the conjecture holds when $M$ is the oriented graphic matroid of a directed graph (Theorem 2.2). In fact, we prove a slightly stronger statement: there exists an exchange sequence of length at most $r$ which is monotone in the sense that the intersection with the target basis does not decrease at every step. We also highlight the connection between our problem and certain conjectures on exchange distances of basis-pairs, and show that no analogous strengthening of the conjecture is possible (Example 3.2).
2 Directed graphs
@@
3 Matroid exchange analogues
@@
There is a clear similarity between Conjectures 1.2 and 3.1, which raises the question of whether the former can be strengthened by allowing symmetric exchanges. More formally, given two $e$-embracing bases $A$ and $B$ of a rank-$r$ oriented matroid, one may ask whether there always exists a sequence of at most $r$ symmetric exchanges transforming $(A,B)$ into $(B,A)$ such that every intermediate pair in the sequence consists of $e$-embracing bases. We show that such a strengthening is not possible, even when the oriented matroid in question is an affine matroid and the bases are disjoint.
@@ Acknowledgement.
Kristóf Bérczi is grateful to Marco Caoduro for introducing him to the original problem on 0-embracing sets, to András Imolay, Benjamin Schröter, and Lorenzo Vecchi for initial discussions on the matroidal extension, and to Frédéric Meunier for pointing out a connection with Hirsch’s conjecture. The research was supported by the Lendület Programme of the Hungarian Academy of Sciences – grant number LP2021-1/2021, by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund – grant numbers ADVANCED 150556 and ELTE TKP 2021-NKTA-62, and by Dynasnet European Research Council Synergy project – grant number ERC-2018-SYG 810115.